My dad is in a blimp. Dad says that when f(x) = sin(x/3.14) is positive .. he is elevated. When f(x) = sin(x/3.14) is negative ... he is below his natural range. Of course, x is in seconds. How many bobs does he take if he adventures for 20 seconds?
Answer by stanbon(75887)

My dad is in a blimp. Dad says that when f(x) = sin(x/3.14) is positive .. he is elevated. When f(x) = sin(x/3.14) is negative ... he is below his natural range. Of course, x is in seconds. How many bobs does he take if he adventures for 20 seconds?
------------
y=sin[(1/pi))x
Period = [2pi/[1/pi] = 2pi^2 =appox 19.739
---------
Assuming your "20 seconds" corresponds to radiuan, the blimp
goes thru one period of a sin curve. You will have to determine
how many "bobs" that is.
===============
cheers,
Stan H.